OpenStudy (anonymous):

A 2-lb bag of brown rice costs five dollars. How much would you expect to pay for a ten-lb bag of brown rice, assuming there is no discount for buying a larger bag?

OpenStudy (anonymous):

@smartone

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

hi, i really need help

OpenStudy (phi):

do you know about ratios or proportions?

OpenStudy (anonymous):

yes

OpenStudy (phi):

write down the ratio of $/pounds for the one you know then write down the ratio for the one you don't know (use x for the part you don't know) and set them equal can you do that ?

OpenStudy (anonymous):

the ratio of dollars to pounds is 5 to 2

OpenStudy (anonymous):

or 5:2

OpenStudy (anonymous):

or 5/2

OpenStudy (phi):

yes , write that as a fraction yes 5/2 = the other ratio

OpenStudy (anonymous):

ok

OpenStudy (phi):

can you write the other ratio? (use x for the unknown cost)

OpenStudy (anonymous):

soooo... ummm 5/2= x ?? sorry i really suck at math

OpenStudy (phi):

the other ratio is also dollars/pounds you know the pounds, right? but you don't know the dollars (so use x for the dollars) can you write the second ratio?

OpenStudy (anonymous):

ohhhh so would it be 10 to x ?

OpenStudy (anonymous):

or 10/x

OpenStudy (anonymous):

or 10:x

OpenStudy (phi):

you want the *same* order as the first ratio. x dollars/ 10 pounds to match 5 dollars/ 2 pounds

OpenStudy (anonymous):

oh my bad so x/10

OpenStudy (phi):

yes, now set the ratios equal to each other

OpenStudy (phi):

you should get 5/2 = x/10

OpenStudy (anonymous):

yes i was just typing that lol

OpenStudy (phi):

now to find x, the fastest way is to multiply both sides by 10 \[ \frac{5}{2} \cdot 10 = \frac{x}{10} \cdot 10 \] notice on the right side, that is the same as \[ \frac{5}{2} \cdot 10 = x\cdot\frac{10}{10} \] and 10/10 is 1 and 1*x is just x, so you get \[ \frac{5}{2} \cdot 10 = x \]

OpenStudy (anonymous):

oh ok

OpenStudy (phi):

can you find what x is ?

OpenStudy (anonymous):

oh duuh yeah

OpenStudy (anonymous):

its 25

OpenStudy (phi):

get the idea? write down the ratio of the part you know write down the ratio (in the *same order* e.g. dollars/pound in this case) use "x" for the part of the ratio you don't know. set the ratios equal to each other solve for x

OpenStudy (anonymous):

oh yeah i do ! thanks so much for your help:D I understand it now!

OpenStudy (phi):

btw, if you end up writing 2/5 = 10/x (which is ok, this is pounds/$ = pounds/$ ) to solve for x, you can first "flip" both sides to get 5/2 = x/10 now solve for x by multiplying both sides by 10

OpenStudy (anonymous):

oh ok i understand. :) thx again @phi